Write the equation of a line parallel to the given line but passing through the given point.
y = x+4; (-7,1)

Answer by Alan3354(69443)

Write the equation of a line parallel to the given line but passing through the given point.
y = x+4; (-7,1)
======================
m = 1
Then do it like this one.
----------------
Slope 5 and the point (4,0)
Use y = mx + b and the point to find b
0 = 5*4 + b
b = -20
-----
y = 5x - 20